OpenStudy (anonymous):

If a person draws 2 cards at random from a standard deck of 52 cards, what is the probability that they get two aces?

OpenStudy (anonymous):

no replacement after each draw?

OpenStudy (anonymous):

assuming no replacement: (4/52)(3/51)=12/2652 =1/221

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@eseidl is this correct 4C2 / 52C2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Yahoo! yes, that works as well